6. A football player makes the opening kickoff and kicks the football with an initial velocity of 28 m/s at an angle of 40° above horizontal. Determine the range of the football.


1
Expert's answer
2021-07-12T12:16:06-0400
"R=\frac{v^2\sin{2\theta}}{g}\\
R=\frac{28^2\sin{80}}{9.8}=78.8\ m"